Hyperspaces of a Moore space and a d-paracompact space Takemi Mizokami, Joetsu, Japan

Keywords: hyperspace, Moore space, d-paracompact AMS Subject classification: 54B20, 54E30 Abstract: In this paper, we investigate the heredity of being a Moore space and being d-paracompact of a space X to its hyperspace K(X) and F(X) with the finite topology.